OpenStudy (anonymous):

Find the linear approximation to f(x) at x = x o. Graph the function and its linear approximation. f(x) = sqrt(x), x o = 1

OpenStudy (anonymous):

@zepdrix can u help?

OpenStudy (amistre64):

linear approx is: \[y=f'(c)(x-c)+f(c)\]

OpenStudy (anonymous):

i tried but i got stuck is this correct so far? L(x) = f(1) + f'(1) (x - 1) = 0 + 1/2sqrt(x) ( x - 1)

OpenStudy (amistre64):

f(1) = 1, not 0 f'(1) = 1/2

OpenStudy (amistre64):

therefore: the linear approximation if say: sqrt(1.2) is\[y=1+\frac12(1.2-1)=1.1\]

OpenStudy (anonymous):

umm so i made a mistake on the 0 right?

OpenStudy (amistre64):

yep, and also on the f'(1) f'(1) not equal 1/2sqrt(x)

OpenStudy (anonymous):

ohh okay

OpenStudy (amistre64):

f'(1) = 1/2sqrt(1) = 1/2

OpenStudy (anonymous):

so its 1 + 0 ( x - 1) ?

OpenStudy (anonymous):

??? is it correct?

OpenStudy (amistre64):

no, 0 does not equal 1/2

OpenStudy (amistre64):

\[y=f(1)+f'(1)(x-1)\] \[y=1+\frac12(x-1)\]

OpenStudy (anonymous):

ohh ok